how many grams of k2cr207are therein 8.90×1024 molecules of k2cr207

1
Expert's answer
2022-03-28T17:06:05-0400

1 mole of matter contains 6.022 × 10²³ units of that substance.

n = m / M

M (K2Cr2O7) = 294.2 g/mol

n (K2Cr2O7) = 8.90×1024 / 6.022 × 10²³ = 14.8 mol

m (K2Cr2O7) = n x M = 14.8 x 294.2 = 4348 g